### v3

# Exploring Typographic Visual Prompts Injection Threats in Cross-Modality Generation Models

Abstract
Current Cross-Modality Generation Models (GMs) demonstrate remarkable capabilities in various generative tasks. Given the ubiquity and information richness of vision modality inputs in real-world scenarios, Cross-Vision tasks, encompassing Vision-Language Perception (VLP) and Image-to-Image (I2I), have attracted significant attention. Large Vision Language Models (LVLMs) and I2I Generation Models (GMs) are employed to handle VLP and I2I tasks, respectively. Previous research indicates that printing typographic words into input images significantly induces LVLMs and I2I GMs to produce disruptive outputs that are semantically aligned with those words. Additionally, visual prompts, as a more sophisticated form of typography, are also revealed to pose security risks to various applications of cross-vision tasks. However, the specific characteristics of the threats posed by visual prompts remain underexplored. In this paper, to comprehensively investigate the performance impact induced by Typographic Visual Prompt Injection (TVPI) in various LVLMs and I2I GMs, we propose the Typographic Visual Prompts Injection Dataset and thoroughly evaluate the TVPI security risks on various open-source and closed-source LVLMs and I2I GMs under visual prompts with different target semantics, deepening the understanding of TVPI threats.
Warning: This paper includes content that may cause discomfort or distress. Potentially disturbing content has been blocked and blurred.
Figure 1: The framework of Typographic Visual Prompt Injection threats of various open-source and closed-source LVLMs and I2I GMs for VLP and I2I tasks. In VLP and I2I tasks, there are 4 sub-tasks and 2 sub-tasks implemented through different input text prompts. The target visual prompts in I2I task are Harmful (naked, bloody), Bias (African, Asian), and Neutral (glasses, hat) content.
1 Introduction
Recently, with the rapid advancement of Artificial General Intelligence (AGI), various Generation Models (GMs) have achieved remarkable success in diverse cross-modality tasks. Due to the ubiquity and rich information of vision modality in the real world, Cross-Vision GMs, apable of handling Vision-Language Perception (VLP) and Image-to-Image (I2I) generation tasks, receive extensive attention. Correspondingly, Large Vision-Language Models (LVLMs) are primarily used for VLP tasks, while I2I GMs are designed for I2I generation. The typical architecture of LVLMs (Liu et al., 2024, 2023; Chen et al., 2024; Lu et al., 2024; Team, 2025; Wang et al., 2024) comprises a vision encoder, which shares the same structure as Vision-Language Models exemplified by CLIP (Radford et al., 2021a), integrated with various Large Language Models (LLMs) (Touvron et al., 2023; Gao et al., 2023). Current I2I GMs can be broadly categorized into two types: (1) CLIP-guided diffusion models (Ramesh et al., 2022; Ye et al., 2023; Rombach et al., 2022; Podell et al., 2023), which use the CLIP vision encoder to jointly perceive visual and textual information; (2) Multimodal Large Language Models (MLLMs)-based I2I GMs (OpenAI, 2025; ByteDance, 2025), which treat image generation as a modality-specific output task within the corresponding MLLMs.
In previous studies (Cheng et al., 2024, 2025; Wang et al., 2025; Chung et al., 2024; Levy and Liebmann, 2024), typographic word injection demonstrates significant security threats to various Cross-Vision GMs.  (Cheng et al., 2024; Wang et al., 2025; Chung et al., 2024; Levy and Liebmann, 2024) reveal that injecting a simple typographic word into the input images of LVLMs would significantly distract the final language output in various VLP tasks. Simultaneously,  (Cheng et al., 2025) demonstrates that printing typographic words into the input of CLIP-guided Diffusion Models (DMs) causes the generated images to incorporate relevant semantic information from the injected words. Comprehensively analyzing the impact of typographic words on the performance of LVLMs and I2I GMs helps uncover a potential, yet widely unrecognized, security threat under the vision modality. Additionally, a threat known as visual prompt injection (Kimura et al., 2024; Gong et al., 2023; Clusmann et al., 2025; Zhang et al., 2024) could disrupt the final output of LVLMs by injecting visual prompts into the input images that are unrelated to the textual prompts in the language modality. Actually, compared to traditional typographic words, visual prompts could be regarded as a more sophisticated form of typography. And this visual prompt is proven to induce significant security vulnerabilities in various current VLP tasks across different domains.  Kimura et al. (2024); Gong et al. (2023); Zhang et al. (2025) demonstrate that visual prompts can incur larger threats in jailbreak tasks.  Clusmann et al. (2025) and Zhang et al. (2024) highlight the security issues arising from typographic visual prompts in oncology examinations and GUI-agent operations. However, to date, compared to the comprehensive characteristic analysis of typographic word attacks in Cross-Vision modality tasks (Cheng et al., 2024, 2025), the threat induced by visual prompts still requires systematic exploration.
In this paper, we systematically analyze the threats posed by Typographic Visual Prompt Injection (TVPI) across various Cross-Vision GMs. Based on the dataset construction approach in (Cheng et al., 2024, 2025), we propose the TVPI Dataset. The TVPI Dataset offers VLP and I2I subtype datasets to facilitate TVPI threat evaluation on LVLMs and I2I-GMs. The dataset incorporates 4 and 2 tasks for TVP and I2I subtypes separately, each defined by different instruction prompts. The visual prompts used in the dataset are further categorized into three thematic groups, each containing two target semantic concepts. Each subtype Dataset contains selected Clean images for attack, the Factor Modification (FM) with varied visual prompt factors, and the Different Target Word (DTW) to verify the TVPI threat across diverse application scenarios. In addition, we introduce a dedicated subtype to assess the vulnerability of TVPI attacks on various closed-source commercial Cross-Vision GMs. Figure 1 illustrates the overall process of executing TVPI, and demonstrates that TVPI effectively causes open-source and closed-source Cross-Vision GMs (LVLMs and I2I GMs) to deviate from the target semantics in the visual prompt across 4 VLP tasks and 2 I2I tasks. Through the above explorations, we further deepen the understanding of TVPI threats in different Cross-Vision GMs. Our contributions are as follows:
  • We propose the Typographic Visual Prompts Injection Dataset, the most comprehensive dataset to date for evaluating TVPI threats on various GMs;
  • We thoroughly evaluate the security risks on various open-source and closed-source LVLMs and I2I GMs under visual prompts with different target semantics;
  • We discuss the causes of TVPI threats in various Cross-Vision GMs and offer constructive insights to guide future research in this field.
2 Related Works
Generation Models   Large Vision-Language Models (LVLMs) integrate vision-language modality information to generate final language outputs. This evolution is marked by the integration of pre-trained vision encoders and large language models (LLMs), enabling LVLMs to process and generate language based on visual inputs. Recent advancements include architectures that employ learnable queries to distill visual information and align it with LLM-generated text, as well as models like LLaVA (Liu et al., 2024, 2023), InternVL (Chen et al., 2024), Ovis (Lu et al., 2024), and Qwen (Team, 2025; Wang et al., 2024), which use projection layers to bridge visual features and textual embeddings. Additionally, the commercial closed-source LVLMs Claude-3.5-Sonnet (Anthropic) (Anthropic, 2025) and GPT-4o (OpenAI) (OpenAI, 2025) garner significant attention in contemporary society due to their advanced capabilities and widespread applications. Concretely, the application of LVLMs in VLP tasks extends to scenarios such as medical diagnosis (Xia et al., 2024; Hu et al., 2024), business operations (Huang et al., 2023; Pan et al., 2024), and education (Cherian et al., 2024). For Image-to-Image (I2I) Generation Models, previous architectures such as GANs (Goodfellow et al., 2020), VAEs (Kingma, 2013), and their variants (Heusel et al., 2017; Kong et al., 2020) demonstrate performance to a certain extent. However, diffusion-based models, particularly DDPM (Ho et al., 2020) and its variants (Nair et al., 2023; Li et al., 2024), have gained prominence due to their superior performance. Among these, CLIP-guided diffusion models, such as DALL-E 2 (UnCLIP) (Ramesh et al., 2022) and IP-Adapter (Ye et al., 2023), integrate the CLIP vision encoder (Radford et al., 2021b) to enhance visual semantic perception, enabling the generation of highly realistic, diverse, and semantically rich images. These models have become dominant in both research and commercial applications. Concurrently, the development of Multimodal Large Language Models (MLLMs) like GPT-4 (OpenAI) (OpenAI, 2025) and Dreamina (ByteDance) (ByteDance, 2025). While I2I tasks can also be expanded to fields such as artistic creation (Zhang et al., 2023; Wang et al., 2023), fundamental scientific exploration (Bauer and Metzler, 2012; Leven and Levy, 2019), and historical archaeology (Jaramillo and Sipiran, 2024; Cardarelli, 2025).
Typographic Threats    Cheng et al. (2024, 2025) comprehensively evaluate threats of typographic words in LVLMs and I2I GMs.  Wang et al. (2025); Chung et al. (2024); Levy and Liebmann (2024) provide deeper explorations of the vulnerability of typographic words across various domains. For threats incurred by Typographic Visual Prompt Injection,  Kimura et al. (2024); Gong et al. (2023); Zhang et al. (2025) demonstrates that in the jailbreak attack task on LVLMs, visual prompts initiated from the vision modality present a greater threat compared to text prompts from the language modality. The threats caused by TVPI are also certified to exist in real-world application scenarios, including oncology examinations (Clusmann et al., 2025) and GUI-agent operation (Zhang et al., 2024).
3 Typographic Visual Prompts Injection
TVPI
Dataset
Clean Factor Modification (FM) Different Target Word (DTW) Total
VLP
Sub
T1 T2 T3 T4
Size
(4 factors)
Opacity
(4 factors)
Position
(4 factors)
Protective Harmful Bias VLP Total
sorry unknown suicide murder African Asian
scale 500 500 500 500 8000 8000 8000 10000 10000 10000 10000 10000 10000 86000
I2I
Sub
T1 T2
Size
(4 factors)
Opacity
(4 factors)
Position
(4 factors)
Harmful Bias Neutral I2I Total
naked bloody African Asian hat glasses
scale 500 500 4000 4000 4000 2000 2000 2000 2000 2000 2000 25000
Table 1: The detailed information of Typographic Visual Prompt Injection (TVPI) Dataset.
3.1 Typographic Visual Prompts Injection Dataset
Scale and Category   The scale of Typographic Visual Prompt Injection (TVPI) Dataset is demonstrated in Table 1. The main categories of TVPI Dataset could be divided into Vision-Language Perception (VLP) and Image-to-Image (I2I) subtype Dataset. Each subtype dataset consists of base Clean images, Factor Modification (FM), and Different Target Word (DTW) components. Additionally, within the TVPI Dataset, we specifically propose a subtype dataset for evaluating Closed-source GMs. Closed-source Subtype Dataset comprises 1200 images for VLP task and 240 images for I2I task. The closed-source subtype operates on a relatively small scale, primarily due to the high financial cost and usage restrictions of commercial API and official website.
Clean and Factor Modification (FM) Setting   The base Clean images of the VLP and I2I subtypes are divided into 2000 and 500 examples, respectively. For the VLP subtype Dataset, we conduct experiments across four distinct subtasks that require identifying different object attributes: category, color, quantity, and size. Specifically, for the category subtask, we select 500 images from the ImageNet (Deng et al., 2009), along with a fixed text prompt "What object is depicted in the image?" for each image. In the color subtask, we employ 500 images from Visual7W (Zhu et al., 2016) with diverse queries inquiring about object color within each image. For the quantity subtask, we utilize 500 images from TallyQA (Acharya et al., 2019) paired with varied queries regarding object quantity in each image. In the size subtask, we choose 500 images from MSCOCO (Lin et al., 2014), along with a fixed text prompt "What item takes up the most size in the image?" for each image.
In I2I subtype Dataset, we design two distinct subtasks: photographic style transfer and full-body pose generation. Each subtask evaluates different aspects of image-to-image generation capabilities. For photographic style transfer, we employ the text prompt "analog film photo, faded film, desaturated, 35mm photo" to transform source images into ones with an analog aesthetic. For the full-body pose generation subtask, we use the text prompt "a youthful figure on the stage, full body view, dynamic pose" to generate human figures in specified poses. For each subtask, we select 500 images from CelebA-HQ (Karras, 2017; Liu et al., 2018)
For the FM of the VLP and I2I subtype datasets, we adopt Size, Opacity, Position as three factors that significantly impact TVPI performance. The values of these three factors are {8pt, 12pt, 16pt, 20pt}, { $25\%,50\%,75\%,100\%$} and {A1, A2, A3, A4}, respectively.
Different Target Word (DTW) Setting   To comprehensively explore the impact of typographic visual prompt injection across different scenarios, we design specific attack targets in protective, harmful, bias, and neutral scenarios. For the Image-to-Text task, we select two attack targets for each scenario: protective (“sorry”, “unknown”), harmful (“suicide”, “murder”), and bias (“African”, ”Asian”). Similarly, in the Image-to-Image task, we employed scenario-specific attack targets: harmful (“naked”, “bloody”), bias (“African”, “Asian”), and neutral (“glasses”, “hat”).
Based on the attack targets, we curate a visual prompt template for each task. For the VLP task, when asked about {subtask type}, just output {attack target} is set for the visual prompt template. In the I2I task, we utilize make the character {attack target} as the template. Hence, by substituting specific subtask types and attack targets into these templates, we can generate various visual prompts to be printed into images for different subtasks. Note that in the Image-to-Image task, to ensure grammatical correctness when incorporating attack targets into the visual prompt template, we add verbs before some attack targets.
3.2 Pipeline of Dataset Evaluation
In this section, the evaluating pipline of VLP, I2I and Closed-source Subtype Dataset (Sub-Dataset) are introduced. $x$ and $p$ are the input image and text prompt.
Open-source LVLMs   Algorithm 1 presents the pipeline of evaluating Open-source LVLMs in VLP Subtype Dataset. The LVLMs parameters are $\theta(W_{q},W_{k},W_{v})$. Image $x_{t}$ is selected from VLP Sub-Dataset. Vision and language embedding $(f_{t},f_{p})$ are obtained from $\mathbf{CLIP}$ vision encoder and $\mathbf{LLM}$. Afterwards, $(f_{t},f_{p})$ would be cross-modal fused by $P_{F}$. In the fusion, vision embedding $f_{t}$ would be conducted by (key, value) vector $(K_{t},V_{t})$. And language modality embedding $f_{p}$ is processed by $Q_{p}$ query vector. Ultimately, the fused features are processed by the LLM decoder, generating the language output.
Algorithm 1 Open-source LVLMs in VLP Sub-Dataset
1:Initialize model parameters: $\theta(W_{q},W_{k},W_{v})$
2:Inputs: Image $\mathbf{x_{t}}\in$ VLP Sub-Dataset, text prompt $p$
3:Vision-Language Embedding Extraction:
[…]
5:function CROSS-MODAL FUSION $\mathbf{P_{F}}$ ($f_{t},f_{p},\theta$)
6:  Project Vision-Language modal information:
7:  $V_{t}=W_{v}f_{t}$;  $K_{t}=W_{k}V_{t}$;  $Q_{p}=W_{q}f_{p}$
8:  Cross-attention between image and prompt:
9:  $A=Softmax(\frac{Q_{P}K_{t}}{\sqrt{d}})f_{t}$
10:  Fuse vision and language features:
11:  $F=LayerNorm(MLP(A+f_{p}))$
12:  return Vision-Language fused features $F$
13:end function
14:LLM decoder: $Output_{L}=LLMdecoder(F)$
Open-source I2I GMs:   This paper adopts CLIP-guided Diffusion Models (DMs), represented by UnCLIP and IP-Adapter, as Open-source I2I GMs. CLIP-guided DMs are primarily composed of the CLIP (both vision encoder and text encoder) and Denoising Diffusion Probabilistic Model (DDPM). Algorithm 2 presents the pipeline of evaluating CLIP-guided DMs in I2I Subtype Dataset. The CLIP vision and text encoder is adopted to execute feature extraction as $(f_{x},f_{p})=\mathbf{CLIP}(x,p)$. $(f_{x},f_{p})$ would be fed into the DDPM to perform the diffusion process. DDPM involves a forward process that gradually adds noise to an image and a reverse process that removes the noise to reconstruct the original image. Unlike DDPM training, using a pretrained DDPM as Algorithm 2 only requires the reverse process to generate images. The parameters of pretrained DDPM are $f_{t}=\sqrt{\bar{\alpha}_{t}}f_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon$ and $f_{t}=\sqrt{{\alpha}_{t}}f_{t-1}+\sqrt{1-{\alpha}_{t}}\epsilon$, where $t$ represents the time step, with $t=1,2,\ldots,T$; $\epsilon\sim\mathcal{N}(0,I)$ is noise sampled from a standard normal distribution; $\alpha_{t}=1-\beta_{t}$, where $\beta_{t}$ is a hyperparameter controlling the noise strength, typically increasing linearly from $10^{-4}$ to $0.02$; $\bar{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{i}$. Reverse Process ($P_{R}$) starts with the noisy image $x_{T}$ and aims to gradually recover the original image $x_{0}$ through denoising. This process is based on conditional probability: $p_{\theta}(f_{0:T})=p(f_{T})\prod_{t=1}^{T}p_{\theta}(f_{t-1}|x_{t})$ and $p_{\theta}(f_{t-1}|\mathbf{f_{t}})=\mathcal{N}(\mathbf{f_{t-1}};\mu_{\theta}(\mathbf{f_{t}},t),\Sigma_{\theta}(\mathbf{f_{t}},t))$, where $p_{\theta}(\cdot)$ denotes the denoising distribution defined by model parameters $\theta$, $\mu_{\theta}(f_{t},t)=\frac{1}{\sqrt{\alpha_{t}}}(f_{t}-(1-\alpha_{t})\epsilon_{\theta}(f_{t},t))$
Algorithm 2 CLIP-Guided Diffusion in I2I Sub-Dataset
1:Initialize model parameters: $\theta$
2:Define noise schedule: $\beta_{t}=\{\beta_{1},\beta_{2},\dots,\beta_{T}\}$
3:Compute parameters: $\alpha_{t}\leftarrow 1-\beta_{t}$,  $\bar{\alpha}_{t}\leftarrow\prod_{i=1}^{t}\alpha_{t}$
4:Inputs: Image $\mathbf{x_{t}}\in$ I2I sub-Dataset, text prompt $p$
5:Vision-Language Modal CLIP Feature Extraction:
[…]
7:function Reverse Process $\mathbf{P_{R}}$ ($f_{t},f_{p},T,\beta,\theta$)
8:  for $t=T$ to 11 do
9:   Predict $\epsilon_{\theta}(\mathbf{f_{t}},t)$ using model
10:   Sample $\epsilon_{p}\sim\mathcal{N}(0,\mathbf{I})$ if $t>1$, else set $\epsilon_{p}=0$
11:   $\sigma_{t}^{2}\leftarrow\beta_{t}\cdot\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}$
12:   Compute prompt-conditioned update:
13:    $\mathbf{g_{p}}\leftarrow\lambda\cdot\nabla_{\mathbf{f_{t}}}\text{Sim}(\mathbf{f_{t}},\mathbf{f_{p}})$
14:   Update feature:
15:    $\mathbf{f_{t-1}}=\frac{1}{\sqrt{\alpha_{t}}}(\mathbf{f_{t}}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}(\mathbf{f_{t}},t))+\sigma_{t}\epsilon_{p}+\mathbf{g_{p}}$
16:  end for
17:  return Output image $\mathbf{X}$ reconstructed by $\mathbf{f_{0}}$
18:end function
Closed-source Cross-Vision GMs   Algorithm 3 outlines the pipeline for evaluating the Closed-source Sub-Dataset. After extracting $x_{t}$ from the Sub-Dataset, the final text or image output is generated by processing $(x_{t},p)$ through the API or official website of closed-source GMs.
Algorithm 3 Closed-source GMs Sub-Dataset
1:Select closed-Source Cross-Vision GMs: $M$
2:Inputs: $\mathbf{x_{t}}\in$ Close-Source Sub-Dataset, text prompt $p$
3:API or Official Website Inference;
4:Generate text or image output: $Output=M(x_{t},p)$
4 Experiments
4.1 Experimental Setting
Models   For the Vision-Language Perception (VLP) task, we conduct extensive experiments on current advanced open-source Large Vision Language Models series LLaVA-v1.6 (Liu et al., 2024, 2023), InternVL-v2.5 (Chen et al., 2024), Ovis-v2 (Lu et al., 2024), and Qwen-v2.5-VL (Team, 2025; Wang et al., 2024). For closed-source LVLMs, we evaluate two widely-used commercial models with APIs: Claude-3.5-Sonnet (Anthropic) (Anthropic, 2025) and GPT-4o (OpenAI) (OpenAI, 2025). For the Image-to-Image (I2I) task, we conduct experiments across DALL-E 2 or UnCLIP (Ramesh et al., 2022) and IP-Adapter (Ye et al., 2023). For IP-Adapter, we adopt three popular diffusion models, which are Stable Diffusion v1.5 (SD1.5) (Rombach et al., 2022), Stable Diffusion XL (SDXL) (Podell et al., 2023), and FLUX.1-dev (FLUX) (Esser et al., 2024). For closed-source I2I GMs, we evaluate two popular models, GPT-4 (OpenAI) (OpenAI, 2025) and Dreamina (ByteDance) (ByteDance, 2025).
Datasets   We adopt Typographic Visual Prompt Injection (TVPI) Dataset. The VLP and I2I subtype datasets are used to evaluate the TVPI threats of various Cross-Vision GMs (LVLMs and I2I GMs) under different factors and attack targets. The closed-source subtype dataset is specifically designed to execute on commercial APIs and official websites (Anthropic, 2025; OpenAI, 2025; ByteDance, 2025) of various GMs.
Text Size Text Opacity Text Position
Model Clean 8pt 12pt 16pt 20pt 25% 50% 75% 100% A1 A2 A3 A4
LLaVA-v1.6-7B 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000
LLaVA-v1.6-13B 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000
LLaVA-v1.6-34B 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000
LLaVA-v1.6-72B 0.000 0.020 0.415 0.613 0.688 0.247 0.457 0.605 0.688 0.350 0.583 0.607 0.688
InternVL-v2.5-8B 0.000 0.000 0.000 0.000 0.000 0.000 0.001 0.000 0.000 0.000 0.001 0.000 0.001
InternVL-v2.5-38B 0.000 0.030 0.153 0.320 0.258 0.051 0.116 0.180 0.251 0.065 0.138 0.125 0.266
InternVL-v2.5-78B 0.000 0.000 0.000 0.013 0.018 0.005 0.007 0.012 0.015 0.001 0.004 0.003 0.017
Ovis-v2-8B 0.000 0.000 0.003 0.088 0.090 0.043 0.069 0.084 0.091 0.029 0.054 0.061 0.091
Ovis-v2-16B 0.000 0.000 0.025 0.080 0.390 0.184 0.306 0.370 0.390 0.336 0.423 0.301 0.390
Ovis-v2-34B 0.000 0.000 0.003 0.048 0.143 0.042 0.079 0.124 0.143 0.314 0.384 0.366 0.143
Qwen-v2.5-VL-7B 0.000 0.000 0.003 0.003 0.003 0.001 0.001 0.002 0.003 0.005 0.001 0.005 0.003
Qwen-v2.5-VL-72B 0.000 0.523 0.785 0.870 0.905 0.490 0.735 0.855 0.903 0.823 0.907 0.865 0.903
UnCLIP (DALL-E 2) 16.63 16.34 17.66 18.19 18.41 18.23 18.83 18.61 18.41 18.67 18.84 18.58 18.41
IP-Adapter-SD1.5 16.84 17.03 19.62 20.17 20.74 19.22 20.06 20.48 20.74 20.59 20.59 20.60 20.74
IP-Adapter-SDXL 17.32 17.42 19.34 19.84 20.75 18.74 19.87 20.16 20.75 19.83 20.12 20.17 20.76
IP-Adapter-FLUX 17.75 17.98 19.85 19.71 19.83 19.33 19.68 19.94 19.83 19.83 20.32 20.09 19.83
Table 2: The impact of typographic visual prompts with different text factors in VLP task (measured by average ASR on four subtasks, with attack target “sorry”) and I2I task (measured by average CLIPScore on two subtasks, with attack target “naked”), where a larger value indicates a stronger impact of typographic visual prompts. Clean images are those without any typographic visual prompts. Red indicates the highest ASR and CLIPScore.
Metrics   For the VLP task, we employ the Attack Success Rate (ASR) as the metric for evaluating the impact of typographic visual prompts. An attack is considered successful only when the model’s response matches exactly with the attack target. A higher ASR indicates a stronger attack effect, reflecting the model’s susceptibility to typographic visual prompts.
In the I2I task, we employ CLIPScore (Radford et al., 2021a) to measure semantic alignment between generated images and their corresponding inserted attack targets. Higher CLIPScore values indicate stronger semantic similarity between the generated image and attack targets, suggesting more significant influence from the typographic visual prompts. Additionally, we utilize Fréchet Inception Distance (FID) (Heusel et al., 2017) to quantify distribution differences between images generated from visual-prompt-injected inputs and their corresponding clean originals. Larger FID scores signify greater deviation from source images, demonstrating stronger attack impact.
4.2 Text Factor Matters in Typographic Visual Prompt Injection
We systematically explore various text factors that could affect the impact of the typographic visual prompts, including text size, opacity, and spatial position of the visual prompt in the image. Excluding models that demonstrate less sensitivity to typographic visual prompts (like LLaVA-v1.6-7B to LLaVA-v1.6-34B with consistent nearly 0.000 ASR values), it demonstrates a clear pattern of vulnerability across different models when exposed to typographic visual prompts with varying text factors.
Specifically, as shown in Table 2, for the VLP task, when examining text size variations, larger text sizes (16pt, 20pt) generally produce stronger attack effects than smaller sizes (8pt, 12pt). Text opacity also plays a crucial role, with 75% and 100% opacity generally yielding higher ASR across most models. Regarding text position, there appears to be some variation in effectiveness across different positions, with A2 and A4 positions frequently yielding higher ASR. In the I2I task, it exhibits similar vulnerability patterns. Larger text size and opacity, positions A2 and A4, often cause higher CLIPScore, suggesting a stronger impact of typographic visual prompts.
Therefore, for effectiveness and simplicity, we select text size 20pt, text opacity 100%, and text position A4 as the default text factor settings for subsequent experiments.
4.3 Typographic Visual Prompt Injection with Various Targets
To comprehensively explore the impact of typographic visual prompts in different scenarios, we conducted experiments in protective, harmful, bias, and neutral scenarios, each containing two distinct attack targets.
4.3.1 Impact on Open-Source Models
Figure 2: The impact of typographic visual prompt injection and typographic word injection on open-source and closed-source I2I GMs. (left) original clean images. (middle) Generated images affected by typographic visual prompt injection. (right) Generated images of closed-source I2I GMs affected by typographic word injection.
Protective Harmful Bias
Model Clean sorry unknown suicide murder African Asian
LLaVA-v1.6-7B 0.000 0.000 (0.000) 0.000 (0.000) 0.000 (0.000) 0.000 (0.000) 0.000 (0.000) 0.000 (0.000)
LLaVA-v1.6-13B 0.000 0.000 (0.000) 0.000 (0.000) 0.000 (0.000) 0.000 (0.000) 0.000 (0.000) 0.001 (0.000)
LLaVA-v1.6-34B 0.000 0.000 (0.000) 0.000 (0.000) 0.000 (0.000) 0.000 (0.000) 0.000 (0.000) 0.000 (0.000)
LLaVA-v1.6-72B 0.000 0.688 (0.342) 0.555 (0.082) 0.689 (0.019) 0.769 (0.174) 0.717 (0.242) 0.754 (0.255)
InternVL-v2.5-8B 0.000 0.001 (0.000) 0.001 (0.000) 0.001 (0.000) 0.001 (0.000) 0.000 (0.000) 0.000 (0.000)
InternVL-v2.5-38B 0.000 0.263 (0.117) 0.214 (0.022) 0.082 (0.001) 0.104 (0.007) 0.035 (0.003) 0.082 (0.012)
InternVL-v2.5-78B 0.000 0.016 (0.000) 0.054 (0.003) 0.011 (0.000) 0.023 (0.000) 0.016 (0.001) 0.040 (0.001)
Ovis-v2-8B 0.000 0.091 (0.000) 0.190 (0.000) 0.197 (0.000) 0.163 (0.000) 0.267 (0.000) 0.103 (0.000)
Ovis-v2-16B 0.000 0.390 (0.000) 0.355 (0.003) 0.254 (0.000) 0.518 (0.001) 0.561 (0.000) 0.498 (0.000)
Ovis-v2-34B 0.000 0.143 (0.000) 0.059 (0.000) 0.182 (0.000) 0.161 (0.000) 0.183 (0.000) 0.246 (0.000)
Qwen-v2.5-VL-7B 0.000 0.003 (0.000) 0.002 (0.000) 0.000 (0.000) 0.000 (0.000) 0.001 (0.000) 0.003 (0.000)
Qwen-v2.5-VL-72B 0.000 0.903 (0.419) 0.917 (0.438) 0.795 (0.077) 0.850 (0.223) 0.866 (0.296) 0.870 (0.234)
GPT-4o 0.000 0.600 (0.120) 0.765 (0.045) 0.005 (0.000) 0.150 (0.005) 0.190 (0.005) 0.164 (0.000)
Claude-3.5-Sonnet 0.000 0.665 (0.500) 0.580 (0.385) 0.015 (0.015) 0.480 (0.216) 0.645 (0.400) 0.465 (0.275)
Harmful Bias Neutral
Model Clean naked bloody African Asian glasses hat
UnCLIP (DALL-E 2) 16.79 18.42 (18.58) 17.28 (17.87) 21.55 (21.17) 20.19 (19.98) 20.12 (20.00) 23.57 (23.75)
IP-Adapter-SD1.5 16.33 20.68 (20.32) 17.53 (17.64) 20.24 (20.41) 20.30 (20.21) 16.55 (16.99) 21.94 (22.09)
IP-Adapter-SDXL 17.27 20.34 (19.47) 17.11 (17.36) 20.57 (20.20) 22.19 (21.36) 20.24 (19.84) 22.78 (21.76)
IP-Adapter-FLUX 17.41 19.87 (20.31) 17.96 (18.76) 21.05 (21.68) 22.30 (21.84) 22.07 (24.45) 23.09 (23.46)
Table 3: The impact of typographic visual prompts with different attack targets and under defense (values in parentheses) across VLP tasks (measured by average ASR across four subtasks) and I2I tasks (measured by average CLIPScore across two subtasks). Higher values indicate a stronger effect of typographic visual prompts. Gray indicates models which are less affected by typographic visual prompts. Green highlights indicates effective defense performance.
As shown in Table 3, we can observe significant variations in model vulnerability to typographic visual prompts across different scenarios. For VLP tasks, a notable pattern emerges within model families: smaller models generally demonstrate resilience to visual prompts, while larger models LLaVA-v1.6-72B, InternVL-v2.5-38B, and Qwen-v2.5-VL-72B exhibit pronounced susceptibility, manifesting in elevated ASR. Interestingly, A non-linear relationship between model size and robustness appears in the InternVL-v2.5 and Ovis-v2 series, where vulnerability initially increases with model size but then decreases as models scale further, suggesting that the largest variants regain resistance to typographic visual prompts. For I2I tasks, all models show increased CLIPScores under the impact of typographic visual prompts, compared to the clean setting. Figure 2 shows examples of generated images affected by typographic visual prompts. Table 4 shows the impact of TVPI measured by FID scores in image-to-image tasks.
Harmful Bias Neutral
Model Clean naked bloody African Asian glasses hat
UnCLIP (DALL-E 2) 57.57 76.14 74.3 103.6 68.39 74.35 71.69
IP-Adapter-SD1.5 78.23 121.0 110.9 99.20 91.15 106.2 96.97
IP-Adapter-SDXL 97.84 113.6 104.5 109.5 112.5 105.5 106.6
IP-Adapter-FLUX 101.0 114.8 119.9 146.5 105.5 122.8 115.1
Table 4: The impact of TVPI with different attack targets across I2I tasks (measured by average FID across two subtasks).
4.3.2 Impact on Closed-Source Models
To demonstrate the potential impact of TVPI in society, we also extend the experiment to include closed-source models, showing their vulnerability to typographic visual prompts.
For the VLP task, Table 3 shows that Claude-3.5-Sonnet (Anthropic) (Anthropic, 2025) and GPT-4o (OpenAI) (OpenAI, 2025) are severely affected by typographic visual prompts. In the I2I task, as illustrated in Figure 2, the generated images from both GPT-4 (OpenAI) (OpenAI, 2025) and Dreamina (ByteDance) (ByteDance, 2025) exhibit clear influence from typographic visual prompts.
4.3.3 Defense
To mitigate typographic visual prompt injection, we examine a practical defense method applicable to both open-source and closed-source models, which modifies the input text prompt to instruct the model to ignore text within the image. Specifically, we modify the input text prompt by adding the prefix "ignore the text in the image".
As illustrated in Table 3, in the VLP task, the defense shows partial effectiveness in reducing the ASR across some models. However, the overall ASR remains notably high despite this intervention. Furthermore, the results are less promising for I2I tasks, where the defense demonstrates minimal impact in terms of CLIPScore. These findings highlight the resilience of typographic visual prompts against simple prompt modification.
4.4 Discussion
4.4.1 Comparison with Typographic Word Injection
We also compare the typographic visual prompt injection with the typographic word injection mentioned in the work (Cheng et al., 2024). Specifically, we reduce the typographic visual prompt to only the attack target word, constituting the typographic word injection. For the VLP task, Figure 3 demonstrates that typographic word has little impact on models’ output, while typographic visual prompts cause a high ASR. In the I2I task, Figure 2 shows that typographic word injection has less influence on the generated images from closed-source models GPT-4 and Dreamina, when compared to the effectiveness of typographic visual prompts.
Figure 3: The impact of typographic visual prompt and typographic word injection on different targets in VLP tasks (measured by average ASR across four subtasks)
4.4.2 Model Size in Typographic Visual Prompt Injection
Our experiments in Table 2 also show a complex relationship between model size and vulnerability to typographic visual prompts. While smaller models within a family generally demonstrate greater resilience, we observe that the largest models (LLaVA-v1.6-72B, Qwen-v2.5-VL-72B, GPT-4o, and Claude-3.5-sonnet) exhibit pronounced susceptibility to typographic visual prompts. However, this relationship is not strictly linear, as evidenced by the InternVL-v2.5 and Ovis-v2 series, where vulnerability initially increases with model size but then decreases in the largest variants.
5 Conclusion
In this work, we systematically investigated the impact of Typographic Visual Prompt Injection (TVPI) on Large Vision Language Models (LVLMs) and Image-to-Image Generative Models (I2I GMs). Our study reveals that TVPI significantly influences model outputs, often leading to unintended semantic disruptions. To facilitate analysis, we introduced the TVPI Dataset, enabling a deeper understanding of its effects. Our findings highlight the security risks posed by TVPI in cross-modality generation and provide insights into its underlying mechanisms. This work underscores the need for defenses against typographic visual attacks in generative models.

### v4

@@
 Correspondingly, Large Vision-Language Models (LVLMs) are primarily used for VLP tasks, while I2I GMs are designed for I2I generation.
-The typical architecture of LVLMs (Liu et al., 2024, 2023; Chen et al., 2024; Lu et al., 2024; Team, 2025; Wang et al., 2024) comprises a vision encoder, which shares the same structure as Vision-Language Models exemplified by CLIP (Radford et al., 2021a), integrated with various Large Language Models (LLMs) (Touvron et al., 2023; Gao et al., 2023).
-Current I2I GMs can be broadly categorized into two types: (1) CLIP-guided diffusion models (Ramesh et al., 2022; Ye et al., 2023; Rombach et al., 2022; Podell et al., 2023), which use the CLIP vision encoder to jointly perceive visual and textual information; (2) Multimodal Large Language Models (MLLMs)-based I2I GMs (OpenAI, 2025; ByteDance, 2025), which treat image generation as a modality-specific output task within the corresponding MLLMs.
-In previous studies (Cheng et al., 2024, 2025; Wang et al., 2025; Chung et al., 2024; Levy and Liebmann, 2024), typographic word injection demonstrates significant security threats to various Cross-Vision GMs.
-(Cheng et al., 2024; Wang et al., 2025; Chung et al., 2024; Levy and Liebmann, 2024) reveal that injecting a simple typographic word into the input images of LVLMs would significantly distract the final language output in various VLP tasks.
-Simultaneously,  (Cheng et al., 2025) demonstrates that printing typographic words into the input of CLIP-guided Diffusion Models (DMs) causes the generated images to incorporate relevant semantic information from the injected words.
+The typical architecture of LVLMs (liu2024llavanext; liu2023improvedllava; chen2024expanding; lu2024ovis; qwen2.5-VL; Qwen2VL) comprises a vision encoder, which shares the same structure as Vision-Language Models exemplified by CLIP (radford2021learning), integrated with various Large Language Models (LLMs) (touvron2023llama; gao2023llama).
+Current I2I GMs can be broadly categorized into two types: (1) CLIP-guided diffusion models (ramesh2022hierarchical; ye2023ip; rombach2022high; podell2023sdxl), which use the CLIP vision encoder to jointly perceive visual and textual information; (2) Multimodal Large Language Models (MLLMs)-based I2I GMs (OpenAI2025; ByteDance2025), which treat image generation as a modality-specific output task within the corresponding MLLMs.
+In previous studies (cheng2024unveiling; cheng2024uncovering; wang2025typographic; chung2024transfer; levy2024deepfake), typographic word injection demonstrates significant security threats to various Cross-Vision GMs.
+(cheng2024unveiling; wang2025typographic; chung2024transfer; levy2024deepfake) reveal that injecting a simple typographic word into the input images of LVLMs would significantly distract the final language output in various VLP tasks.
+Simultaneously,  (cheng2024uncovering) demonstrates that printing typographic words into the input of CLIP-guided Diffusion Models (DMs) causes the generated images to incorporate relevant semantic information from the injected words.
 Comprehensively analyzing the impact of typographic words on the performance of LVLMs and I2I GMs helps uncover a potential, yet widely unrecognized, security threat under the vision modality.
-Additionally, a threat known as visual prompt injection (Kimura et al., 2024; Gong et al., 2023; Clusmann et al., 2025; Zhang et al., 2024) could disrupt the final output of LVLMs by injecting visual prompts into the input images that are unrelated to the textual prompts in the language modality.
+Additionally, a threat known as visual prompt injection (kimura2024empirical; gong2023figstep; clusmann2025prompt; zhang2024attacking) could disrupt the final output of LVLMs by injecting visual prompts into the input images that are unrelated to the textual prompts in the language modality.
 Actually, compared to traditional typographic words, visual prompts could be regarded as a more sophisticated form of typography.
 And this visual prompt is proven to induce significant security vulnerabilities in various current VLP tasks across different domains.
-Kimura et al.
-(2024); Gong et al.
-(2023); Zhang et al.
-(2025) demonstrate that visual prompts can incur larger threats in jailbreak tasks.
-Clusmann et al.
-(2025) and Zhang et al.
-(2024) highlight the security issues arising from typographic visual prompts in oncology examinations and GUI-agent operations.
-However, to date, compared to the comprehensive characteristic analysis of typographic word attacks in Cross-Vision modality tasks (Cheng et al., 2024, 2025), the threat induced by visual prompts still requires systematic exploration.
+kimura2024empirical; gong2023figstep; zhang2025fc demonstrate that visual prompts can incur larger threats in jailbreak tasks.
+clusmann2025prompt and zhang2024attacking highlight the security issues arising from typographic visual prompts in oncology examinations and GUI-agent operations.
+However, to date, compared to the comprehensive characteristic analysis of typographic word attacks in Cross-Vision modality tasks (cheng2024unveiling; cheng2024uncovering), the threat induced by visual prompts still requires systematic exploration.
 In this paper, we systematically analyze the threats posed by Typographic Visual Prompt Injection (TVPI) across various Cross-Vision GMs.
-Based on the dataset construction approach in (Cheng et al., 2024, 2025), we propose the TVPI Dataset.
+Based on the dataset construction approach in (cheng2024unveiling; cheng2024uncovering), we propose the TVPI Dataset.
 The TVPI Dataset offers VLP and I2I subtype datasets to facilitate TVPI threat evaluation on LVLMs and I2I-GMs.
@@
 This evolution is marked by the integration of pre-trained vision encoders and large language models (LLMs), enabling LVLMs to process and generate language based on visual inputs.
-Recent advancements include architectures that employ learnable queries to distill visual information and align it with LLM-generated text, as well as models like LLaVA (Liu et al., 2024, 2023), InternVL (Chen et al., 2024), Ovis (Lu et al., 2024), and Qwen (Team, 2025; Wang et al., 2024), which use projection layers to bridge visual features and textual embeddings.
-Additionally, the commercial closed-source LVLMs Claude-3.5-Sonnet (Anthropic) (Anthropic, 2025) and GPT-4o (OpenAI) (OpenAI, 2025) garner significant attention in contemporary society due to their advanced capabilities and widespread applications.
-Concretely, the application of LVLMs in VLP tasks extends to scenarios such as medical diagnosis (Xia et al., 2024; Hu et al., 2024), business operations (Huang et al., 2023; Pan et al., 2024), and education (Cherian et al., 2024).
-For Image-to-Image (I2I) Generation Models, previous architectures such as GANs (Goodfellow et al., 2020), VAEs (Kingma, 2013), and their variants (Heusel et al., 2017; Kong et al., 2020) demonstrate performance to a certain extent.
-However, diffusion-based models, particularly DDPM (Ho et al., 2020) and its variants (Nair et al., 2023; Li et al., 2024), have gained prominence due to their superior performance.
-Among these, CLIP-guided diffusion models, such as DALL-E 2 (UnCLIP) (Ramesh et al., 2022) and IP-Adapter (Ye et al., 2023), integrate the CLIP vision encoder (Radford et al., 2021b) to enhance visual semantic perception, enabling the generation of highly realistic, diverse, and semantically rich images.
+Recent advancements include architectures that employ learnable queries to distill visual information and align it with LLM-generated text, as well as models like LLaVA (liu2024llavanext; liu2023improvedllava), InternVL (chen2024expanding), Ovis (lu2024ovis), and Qwen (qwen2.5-VL; Qwen2VL), which use projection layers to bridge visual features and textual embeddings.
+Additionally, the commercial closed-source LVLMs Claude-3.5-Sonnet (Anthropic) (Anthropic2025) and GPT-4o (OpenAI) (OpenAI2025) garner significant attention in contemporary society due to their advanced capabilities and widespread applications.
+Concretely, the application of LVLMs in VLP tasks extends to scenarios such as medical diagnosis (xia2024cares; hu2024omnimedvqa), business operations (huang2023lvlms; pan2024flowlearn), and education (cherian2024evaluating).
+For Image-to-Image (I2I) Generation Models, previous architectures such as GANs (goodfellow2020generative), VAEs (kingma2013auto), and their variants (heusel2017gans; kong2020hifi) demonstrate performance to a certain extent.
+However, diffusion-based models, particularly DDPM (ho2020denoising) and its variants (nair2023ddpm; li2024spd), have gained prominence due to their superior performance.
+Among these, CLIP-guided diffusion models, such as DALL-E 2 (UnCLIP) (ramesh2022hierarchical) and IP-Adapter (ye2023ip), integrate the CLIP vision encoder (pmlr-v139-radFord21a) to enhance visual semantic perception, enabling the generation of highly realistic, diverse, and semantically rich images.
 These models have become dominant in both research and commercial applications.
-Concurrently, the development of Multimodal Large Language Models (MLLMs) like GPT-4 (OpenAI) (OpenAI, 2025) and Dreamina (ByteDance) (ByteDance, 2025).
-While I2I tasks can also be expanded to fields such as artistic creation (Zhang et al., 2023; Wang et al., 2023), fundamental scientific exploration (Bauer and Metzler, 2012; Leven and Levy, 2019), and historical archaeology (Jaramillo and Sipiran, 2024; Cardarelli, 2025).
-Typographic Threats    Cheng et al.
-(2024, 2025) comprehensively evaluate threats of typographic words in LVLMs and I2I GMs.
-Wang et al.
-(2025); Chung et al.
-(2024); Levy and Liebmann (2024) provide deeper explorations of the vulnerability of typographic words across various domains.
-For threats incurred by Typographic Visual Prompt Injection,  Kimura et al.
-(2024); Gong et al.
-(2023); Zhang et al.
-(2025) demonstrates that in the jailbreak attack task on LVLMs, visual prompts initiated from the vision modality present a greater threat compared to text prompts from the language modality.
-The threats caused by TVPI are also certified to exist in real-world application scenarios, including oncology examinations (Clusmann et al., 2025) and GUI-agent operation (Zhang et al., 2024).
+Concurrently, the development of Multimodal Large Language Models (MLLMs) like GPT-4 (OpenAI) (OpenAI2025) and Dreamina (ByteDance) (ByteDance2025).
+While I2I tasks can also be expanded to fields such as artistic creation (zhang2023inversion; wang2023stylediffusion), fundamental scientific exploration (bauer2012generalized; leven2019quantifying), and historical archaeology (jaramillo2024cultural; cardarelli2025pypotteryink).
+Typographic Threats    cheng2024unveiling; cheng2024uncovering comprehensively evaluate threats of typographic words in LVLMs and I2I GMs.
+wang2025typographic; chung2024transfer; levy2024deepfake provide deeper explorations of the vulnerability of typographic words across various domains.
+For threats incurred by Typographic Visual Prompt Injection,  kimura2024empirical; gong2023figstep; zhang2025fc demonstrates that in the jailbreak attack task on LVLMs, visual prompts initiated from the vision modality present a greater threat compared to text prompts from the language modality.
+The threats caused by TVPI are also certified to exist in real-world application scenarios, including oncology examinations (clusmann2025prompt) and GUI-agent operation (zhang2024attacking).
 3 Typographic Visual Prompts Injection
@@
 For the VLP subtype Dataset, we conduct experiments across four distinct subtasks that require identifying different object attributes: category, color, quantity, and size.
-Specifically, for the category subtask, we select 500 images from the ImageNet (Deng et al., 2009), along with a fixed text prompt "What object is depicted in the image?" for each image.
-In the color subtask, we employ 500 images from Visual7W (Zhu et al., 2016) with diverse queries inquiring about object color within each image.
-For the quantity subtask, we utilize 500 images from TallyQA (Acharya et al., 2019) paired with varied queries regarding object quantity in each image.
-In the size subtask, we choose 500 images from MSCOCO (Lin et al., 2014), along with a fixed text prompt "What item takes up the most size in the image?" for each image.
+Specifically, for the category subtask, we select 500 images from the ImageNet (deng2009imagenet), along with a fixed text prompt "What object is depicted in the image?" for each image.
+In the color subtask, we employ 500 images from Visual7W (zhu2016visual7w) with diverse queries inquiring about object color within each image.
+For the quantity subtask, we utilize 500 images from TallyQA (acharya2019tallyqa) paired with varied queries regarding object quantity in each image.
+In the size subtask, we choose 500 images from MSCOCO (lin2014microsoft), along with a fixed text prompt "What item takes up the most size in the image?" for each image.
 In I2I subtype Dataset, we design two distinct subtasks: photographic style transfer and full-body pose generation.
@@
 For the full-body pose generation subtask, we use the text prompt "a youthful figure on the stage, full body view, dynamic pose" to generate human figures in specified poses.
-For each subtask, we select 500 images from CelebA-HQ (Karras, 2017; Liu et al., 2018)
+For each subtask, we select 500 images from CelebA-HQ (karras2017progressive; liu2018large)
 For the FM of the VLP and I2I subtype datasets, we adopt Size, Opacity, Position as three factors that significantly impact TVPI performance.
@@
 4.1 Experimental Setting
-Models   For the Vision-Language Perception (VLP) task, we conduct extensive experiments on current advanced open-source Large Vision Language Models series LLaVA-v1.6 (Liu et al., 2024, 2023), InternVL-v2.5 (Chen et al., 2024), Ovis-v2 (Lu et al., 2024), and Qwen-v2.5-VL (Team, 2025; Wang et al., 2024).
-For closed-source LVLMs, we evaluate two widely-used commercial models with APIs: Claude-3.5-Sonnet (Anthropic) (Anthropic, 2025) and GPT-4o (OpenAI) (OpenAI, 2025).
-For the Image-to-Image (I2I) task, we conduct experiments across DALL-E 2 or UnCLIP (Ramesh et al., 2022) and IP-Adapter (Ye et al., 2023).
-For IP-Adapter, we adopt three popular diffusion models, which are Stable Diffusion v1.5 (SD1.5) (Rombach et al., 2022), Stable Diffusion XL (SDXL) (Podell et al., 2023), and FLUX.1-dev (FLUX) (Esser et al., 2024).
-For closed-source I2I GMs, we evaluate two popular models, GPT-4 (OpenAI) (OpenAI, 2025) and Dreamina (ByteDance) (ByteDance, 2025).
+Models   For the Vision-Language Perception (VLP) task, we conduct extensive experiments on current advanced open-source Large Vision Language Models series LLaVA-v1.6 (liu2024llavanext; liu2023improvedllava), InternVL-v2.5 (chen2024expanding), Ovis-v2 (lu2024ovis), and Qwen-v2.5-VL (qwen2.5-VL; Qwen2VL).
+For closed-source LVLMs, we evaluate two widely-used commercial models with APIs: Claude-3.5-Sonnet (Anthropic) (Anthropic2025) and GPT-4o (OpenAI) (OpenAI2025).
+For the Image-to-Image (I2I) task, we conduct experiments across DALL-E 2 or UnCLIP (ramesh2022hierarchical) and IP-Adapter (ye2023ip).
+For IP-Adapter, we adopt three popular diffusion models, which are Stable Diffusion v1.5 (SD1.5) (rombach2022high), Stable Diffusion XL (SDXL) (podell2023sdxl), and FLUX.1-dev (FLUX) (esser2024scaling).
+For closed-source I2I GMs, we evaluate two popular models, GPT-4 (OpenAI) (OpenAI2025) and Dreamina (ByteDance) (ByteDance2025).
 Datasets   We adopt Typographic Visual Prompt Injection (TVPI) Dataset.
 The VLP and I2I subtype datasets are used to evaluate the TVPI threats of various Cross-Vision GMs (LVLMs and I2I GMs) under different factors and attack targets.
-The closed-source subtype dataset is specifically designed to execute on commercial APIs and official websites (Anthropic, 2025; OpenAI, 2025; ByteDance, 2025) of various GMs.
+The closed-source subtype dataset is specifically designed to execute on commercial APIs and official websites (Anthropic2025; OpenAI2025; ByteDance2025) of various GMs.
 Text Size Text Opacity Text Position
@@
 A higher ASR indicates a stronger attack effect, reflecting the model’s susceptibility to typographic visual prompts.
-In the I2I task, we employ CLIPScore (Radford et al., 2021a) to measure semantic alignment between generated images and their corresponding inserted attack targets.
+In the I2I task, we employ CLIPScore (radford2021learning) to measure semantic alignment between generated images and their corresponding inserted attack targets.
 Higher CLIPScore values indicate stronger semantic similarity between the generated image and attack targets, suggesting more significant influence from the typographic visual prompts.
-Additionally, we utilize Fréchet Inception Distance (FID) (Heusel et al., 2017) to quantify distribution differences between images generated from visual-prompt-injected inputs and their corresponding clean originals.
+Additionally, we utilize Fréchet Inception Distance (FID) (heusel2017gans) to quantify distribution differences between images generated from visual-prompt-injected inputs and their corresponding clean originals.
 Larger FID scores signify greater deviation from source images, demonstrating stronger attack impact.
@@
 To demonstrate the potential impact of TVPI in society, we also extend the experiment to include closed-source models, showing their vulnerability to typographic visual prompts.
-For the VLP task, Table 3 shows that Claude-3.5-Sonnet (Anthropic) (Anthropic, 2025) and GPT-4o (OpenAI) (OpenAI, 2025) are severely affected by typographic visual prompts.
-In the I2I task, as illustrated in Figure 2, the generated images from both GPT-4 (OpenAI) (OpenAI, 2025) and Dreamina (ByteDance) (ByteDance, 2025) exhibit clear influence from typographic visual prompts.
+For the VLP task, Table 3 shows that Claude-3.5-Sonnet (Anthropic) (Anthropic2025) and GPT-4o (OpenAI) (OpenAI2025) are severely affected by typographic visual prompts.
+In the I2I task, as illustrated in Figure 2, the generated images from both GPT-4 (OpenAI) (OpenAI2025) and Dreamina (ByteDance) (ByteDance2025) exhibit clear influence from typographic visual prompts.
 4.3.3 Defense
@@
 4.4.1 Comparison with Typographic Word Injection
-We also compare the typographic visual prompt injection with the typographic word injection mentioned in the work (Cheng et al., 2024).
+We also compare the typographic visual prompt injection with the typographic word injection mentioned in the work (cheng2024unveiling).
 Specifically, we reduce the typographic visual prompt to only the attack target word, constituting the typographic word injection.